Consider the following probability distribution for X. 7 0.1 0.1 0.2 003 0.2 0.1 (i) (ii) (iii) (iv) Find E(X). Find E(2X3+4x) Determine the MGF of X. Calculate Var (X) using MGF of X.
Answer #1

i)

E(X)=\sumxP(x)=-3*(0.1)+(-1)*0.2+1*0.3+3*0.2+5*0.1+7*0.1=1.6

ii)

E(2x3+4X) =\sum(2x3+4X)*P(x)=(2*(-3)3+4*(-3))*(0.1)+(2*(-1)3+4*(-1))*0.2+(2*(1)3+4*1)*0.3+(2*33+4*3)*0.2+(2*53+4*5)*0.1+(2*73+4*2)*0.1=984

iii)

MGF =Mx(t)=E(etx) =\sumetx*P(x)=0.1*e-3t+0.2*e-t+0.3*et+0.2*e3t+0.1*e5t+0.1*e7t

here first derivation of mgf =Mx'(t)=-0.3*e-3t-0.2*e-t+0.3*et+0.6*e3t+0.5*e5t+0.7*e7t

therefore E(X)=Mx'(0)=-0.3*e-3*0-0.2*e-0+0.3*e0+0.6*e3*0+0.5*e5*0+0.7*e7*0 = 1.6

second derivative of mgf Mx ''(t)=0.9*e-3t+0.2*e-t+0.3*et+1.8*e3t+2.5*e5t+4.9*e7t

E(X2)=Mx ''(0)=0.9*e-3*0+0.2*e-0+0.3*e0+1.8*e3*0+2.5*e5*0+4.9*e7*0 =10.6

hence Var(X)=E(X2)-(E(X))2 =10.6-1.62 =8.04